OpenStudy (anonymous):

elimination 3x+2y=0 and -3x-y=3

OpenStudy (bahrom7893):

from the 2nd equation: y = -3-3x

OpenStudy (bahrom7893):

Pluggin that into first equation: 3x - 6 - 6x = 0 -3x = 6 x = -2 y = -3 - 3*2 = -9